a resterant manager want to purchase 200sets of dishes. One design cost $25 per set,while another cost $45 dollars per set. If she want to use her entire budget of $7800, how many of each design should be ordered?
Answer by stanbon(57361)

a restaurant manager wants to purchase 200 sets of dishes. One design costs $25 per set,while another costs $45 dollars per set. If she wants to use her entire budget of $7800, how many of each design should be ordered?
------------------
t + f = 200 sets
25t+45f = 7800 dollars
-----------------------------------
Modify the 1st equation:
25t + 25f = 25*200
25t + 45f = 7800
-----
Subtract and solve for "f":
20f = 2800
f = 140 (# of $45 dollars sets to purchase)
-----
Solve for "t":
t + f = 200
t + 140 = 200
t = 60 (# of $25 sets to purchase)
=============
Cheers,
Stan H.